# HyperSHAP: Shapley Values and Interactions for Hyperparameter Importance

### v1

@@ 4.1 Ablation of Hyperparameter Configurations
One common scenario for quantifying the HPI is to compare a hyperparameter configuration (HPC) $\bm{\lambda}^{\ast}$ of interest to some reference HPC $\bm{\lambda}^{0}$, e.g., the default parameterization of a learner as provided by its implementing library or a tuned default HPC that has proven effective for past tasks. In turn, $\bm{\lambda}^{\ast}$ can be any HPC obtained through HPO or a manual configuration. Given $\bm{\lambda}^{\ast}$ and $\bm{\lambda}^{0}$, the question now is how values of $\bm{\lambda}^{\ast}$ affect the performance of the learner relative to the reference HPC $\bm{\lambda}^{0}$. To this end, we can transition from the reference HPC to the HPC of interest by switching the values of hyperparameters one by one from its value in $\bm{\lambda}^{0}$ to the value in $\bm{\lambda}^{\ast}$, which is also done in empirical ML studies and referred to as ablations.
Ablation analysis has already been followed by Fawcett & Hoos (2016) and Biedenkapp et al. (2017) but restricted to single hyperparameter ablations that are executed sequentially. Instead, we consider the HPI game of Ablation using all possible subsets, which allows us to capture interactions.
Definition 4.1 (HPI Game - Ablation).
The Ablation HPI game $\nu_{G_{A}}:2^{\mathcal{N}}\to\mathbb{R}$ is defined based on a tuple
$G_{A}:=(\bm{\lambda}^{0},\bm{\lambda}^{\ast},D,u),$
consisting of a baseline (default) HPC $\bm{\lambda}^{0}$, an HPC of interest $\bm{\lambda}^{\ast}$, a dataset $D$, and a measure $u$. Given a coalition $S\subseteq\mathcal{N}$, we construct an intermediate HPC with $\oplus_{S}:\bm{\Lambda}\times\bm{\Lambda}\to\bm{\Lambda}$ as
@@
and evaluate its worth with
$\nu_{G_{A}}(S):=\textsc{Val}_{u}(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0},D)\,\,.$
The Ablation game quantifies the worth of a coalition based on the comparison with a default HPC $\bm{\lambda}^{0}$. In XAI terminology, this approach is known as baseline imputation, cf. Section 3. Natural extensions of the Ablation game capture these ablations with respect to a distribution $\bm{\lambda}^{0}\sim p^{0}(\bm{\lambda}^{0})$ over the baseline HPC space $\bm{\Lambda}$ as
$\mathbb{E}_{\bm{\lambda}^{0}\sim p^{0}(\bm{\lambda}^{0})}[\textsc{Val}_{u}(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0},D)]\,\,,$
which relates to the marginal performance introduced by Hutter et al. (2014). In XAI terminology, it is further distinguished between distributions $p(\bm{\lambda}^{0})$ that either depend (conditional) or do not depend (marginal) on the HPC of interest $\bm{\lambda}^{\ast}$. While baseline imputation is mostly chosen for computational efficiency, it was argued that it also satisfies beneficial properties (Sundararajan & Najmi, 2020). Nevertheless, the choice of a baseline has a strong impact on the explanation (Sturmfels et al., 2020). In HPO, we are typically given a default HPC $\bm{\lambda}^{0}$, which we use for the Ablation game, but our methodology can be directly extended to the probabilistic setting.
4.2 Tunability of Learners
Zooming out from a specific configuration, we can ask to what extent it is worthwhile to tune hyperparameters. In the literature, this question has been connected to the term of tunability (Probst et al., 2019). Tunability aims to quantify how much performance improvements can be obtained by tuning a learner comparing against a baseline HPC, e.g., an HPC that is known to work well across various datasets (Pushak & Hoos, 2020). In this context, we are interested in the importance of tuning specific hyperparameters. A classical tool to quantify variable importance is sensitivity analysis (Owen, 2013), which measures the variance induced by the variables of interest and decomposes their contributions into main and interaction effects.
Definition 4.2 (HPI Game - Sensitivity).
The Sensitivity game $\nu_{G_{V}}:2^{\mathcal{N}}\to\mathbb{R}$ is defined based on a tuple
$G_{V}:=(\bm{\lambda}^{0},\bm{\Lambda},p^{*},D,u),$
consisting of a baseline HPC $\bm{\lambda}^{0}$, an HPC space of interest $\bm{\Lambda}$ equipped with a probability distribution $p^{*}$, a dataset $D$, and a measure $u$. The value function is given by
@@ 4.2 Tunability of Learners
A large value of a coalition $S\subseteq\mathcal{N}$ in the Sensitivity game indicates that these hyperparameters are important to be set to the right value. Hutter et al. (2014) implicitly rely on the Sensitivity game and compute the functional ANOVA decomposition, quantifying pure main and interaction effects. In game theory, this corresponds to the MIs of the Sensitivity game,which can be summarized into interpretable representations using the SV and SIs (Fumagalli et al., 2024a).
While sensitivity analysis is a suitable tool in XAI, it has some drawbacks for Tunability. First, as illustrated below, the total variance being decomposed $\nu_{G_{V}}(\mathcal{N})$ strongly depends on the chosen probability distribution $p^{*}$ and the HPC space $\bm{\Lambda}$. Moreover, it does not reflect the performance increase expected when tuning all hyperparameters. Second, for a coalition of hyperparameters $S\subseteq\mathcal{N}$, we expect that the coalition’s worth (performance) increases when tuning additional hyperparameters, i.e., $\nu(S)\leq\nu(T)$, if $S\subseteq T$. This property is known as monotonicity (Fujimoto et al., 2006), but does not hold in general for the Sensitivity game $\nu_{G_{V}}$. For a simple example, we refer to Section A.3. Instead, we now propose the monotone Tunability HPI game.
Definition 4.3 (HPI Game - Tunability).
The Tunability HPI game is defined by a tuple
$G_{T}=(\bm{\lambda}^{0},\bm{\Lambda},D,u),$
consisting of a baseline HPC $\bm{\lambda}^{0}\in\bm{\Lambda}$, an HPC space $\bm{\Lambda}$, a dataset $D$, and a measure $u$. The value function is given by
@@
The Tunability game directly measures the performance obtained from tuning the hyperparameters of a coalition $S$ while leaving the remaining hyperparameters at the default value $\bm{\lambda}^{0}$. The Tunability game is monotone, which yields the following theorem.
Theorem 4.4.
@@
While the main effects obtained from the Tunability game are non-negative, interactions clearly can be negative, indicating redundancies of the involved hyperparameters.
Benefits of Tunability over Sensitivity.
We now showcase the benefits of the Tunability game over the Sensitivity game using a synthetic example. We consider a two-dimensional HPC space $\bm{\Lambda}:=\Lambda_{1}\times\Lambda_{2}$ with discrete HPCs $\Lambda_{1}:=\{0,1\}$ and $\Lambda_{2}:=\{0,\dots,m\}$ for $m>1$. The optimal configuration is defined as $\bm{\lambda}^{\ast}:=(1,m)$, and the performance is quantified by $\textsc{Val}_{u}(\bm{\lambda},D):=\mathbf{1}_{\lambda_{1}=\lambda_{1}^{\ast}}+\mathbf{1}_{\lambda_{2}=\lambda^{\ast}_{2}}$, where $\mathbf{1}$ is the indicator function. That is, we observe an increase of performance of $1$ for each of the hyperparameters set to the optimal HPC $\bm{\lambda}^{\ast}$. Lastly, we set the HPC baseline to $\bm{\lambda}^{0}:=(0,0)$ or $\bm{\lambda}^{0}:=\bm{\lambda}^{\ast}$. Intuitively, we expect that both hyperparameters obtain similar HPI scores, since they both contribute equally to the optimal performance $\textsc{Val}_{u}(\bm{\lambda}^{\ast},D)=2$. Moreover, if the baseline is set to the optimal HPC $\bm{\lambda}^{\ast}$, we expect the HPI to reflect that there is no benefit of tuning. Since the hyperparameters independently affect the performance, we do not expect any interactions.
Theorem 4.5.
The HPI scores of the Sensitivity and Tunability game for the synthetic example are given by Table 1.
Both HPI scores correctly quantify the absence of interaction $\lambda_{1}\times\lambda_{2}$. In contrast to the Tunability game, the Sensitivity game assigns smaller HPI scores to the hyperparameter $\lambda_{2}$ due to the larger domain $\Lambda_{2}$. In fact, the Sensitivity HPI score of $\lambda_{2}$ roughly decreases with order $m^{-1}$. Moreover, the Tunability HPI scores reflect the performance increase and decompose the difference between the optimal and the default performance. In contrast, the Sensitivity HPI scores decompose the overall variance, which depends on $\bm{\Lambda}$ and $p^{*}$. Lastly, setting the default HPC $\bm{\lambda}^{0}$ to $\bm{\lambda}^{\ast}$ decreases the Tunability HPI scores to zero, whereas the Sensitivity HPI scores remain unaffected. In summary, Sensitivity reflects the variability in performance when changing the hyperparameters, whereas Tunability reflects the benefit of tuning.
Table 1: HPI main and interaction effects of a two-dimensional synthetic HPO problem for the Sensitivity and Tunability game with baseline HPC set to $(0,0)$ and the optimal HPC $\bm{\lambda}^{\ast}$. The Sensitivity game assigns smaller contributions to hyperparameters with a larger domain ($\lambda_{2})$. Setting $\bm{\lambda}^{0}=\bm{\lambda}^{*}$ reduces the Tunability HPI scores to zero, whereas Sensitivity is unaffected.
HPI Game Sensitivity Tunability
@@ 4.3 Optimizer Bias
The Tunability game aims to explain the importance of hyperparameters being tuned, which can also be used to gain insights into the capabilities of a hyperparameter optimizer. In particular, by comparing the optimal performance with the empirical performance of a single optimizer, we uncover biases and point to specific hyperparameters that the optimizer fails to exploit. We define a hyperparameter optimizer as a function $\mathcal{O}:\mathbb{D}\times 2^{\bm{\Lambda}}\rightarrow\bm{\Lambda}$, mapping from the space of datasets and an HPC space to an HPC.
Definition 4.6 (HPI Game - Optimizer Bias).
The Optimizer Bias HPI game is defined as a tuple
$G_{O}=(\bm{\Lambda},\bm{\lambda}^{0},\mathcal{O},D,u),$
consisting of a HPC space $\bm{\Lambda}$, a baseline $\bm{\lambda}^{0}$, the hyperparameter optimizer of interest $\mathcal{O}$, a dataset $D$ and a measure $u$. For $S\subseteq\mathcal{N}$, we construct $\bm{\Lambda}^{S}:=\{\bm{\lambda}\oplus_{S}\bm{\lambda}^{0}:\bm{\lambda}\in\bm{\Lambda}\}$ and define
@@
Intuitively speaking, the value function measures any deviation from the performance of the actual best-performing HPC. In other words, with the help of Definition 4.6, we can identify deficiencies of the hyperparameter optimizer $\mathcal{O}$ over the best-performing solution and, thereby for example, identify whether an optimizer struggles to optimize certain (types of) hyperparameters.
@@ 4.4 Practical Aspects of HyperSHAP
Naively, to evaluate a single coalition in Definition 4.3, we need to conduct one HPO run. While this can be costly, we argue that using surrogate models that are, e.g., obtained through Bayesian optimization, can be used to simulate HPO, rendering HyperSHAP more tractable. This is similar to other surrogate-based explainability methods for HPO (Hutter et al., 2014; Biedenkapp et al., 2017; Moosbauer et al., 2021; Segel et al., 2023). In contrast, to analyze Optimizer Bias, we propose to approximate $\nu_{G_{T}}$ using a diverse ensemble of optimizers $\mathbb{O}:=\{\mathcal{O}_{i}\}$, and choose the best result for $\bm{\Lambda}^{S}$ obtained through any optimizer from $\mathbb{O}$, forming a virtual optimizer that always yields the best-known value. This virtual best hyperparameter optimizer (VBO) approximates
@@
HPI Game Extensions across Multiple Datasets.
In a more general setting, we are also interested in the different aspects of HyperSHAP across multiple datasets, where we present direct extensions of the previous games.
Definition 4.7 (HPI Game- Multi-Dataset Variants).
Given a collection of datasets $\mathcal{D}:=\{D_{1},\dots,D_{M}\}$ and the corresponding HPI games $\nu^{D_{i}}_{G}$ for $i\in\{1,\dots,M\}$ with $G\in\{G_{A},G_{V},G_{T},G_{O}\}$, we define its multi-dataset variant with the value function
$\nu^{\mathcal{D}}_{G}(S):=\bigoplus_{i=1}^{M}\nu^{D_{i}}_{G}(S),$
where $\bigoplus$ denotes an aggregation operator, e.g. the mean, of the game values obtained from the individual datasets $D_{i}$.
Considering HPI across datasets allows for a broader and more comprehensive assessment of the impact of individual hyperparameters and how they interact with each other. By aggregating the value of a coalition over datasets, we can evaluate the generalizability of HPI in contrast to identifying which hyperparameter is important for which dataset. For instance, this can justify recommendations with respect to Multi-Data Tunability, which hyperparameters should be tuned in general. Alternatively, we can assure a systematic optimizer bias with respect to certain hyperparameters instead of observing data-specific effects only.
Figure 2: Upset plots (Lex et al., 2014) showing the hyperparameter importances for Ablation (left) and Tunability (right) wrt. lm1b_transformer from PD1 (Wang et al., 2024).

### v2

@@ 5.1 Ablation of Hyperparameter Configurations
One common approach to explaining HPO results is to compare a configuration of interest, $\bm{\lambda}^{\ast}$, to a reference configuration $\bm{\lambda}^{0}$, typically a library default or a tuned default that has performed well on prior tasks. The configuration $\bm{\lambda}^{\ast}$ may stem from HPO or be manually selected. The key question is how changes in $\bm{\lambda}^{\ast}$ impact performance relative to $\bm{\lambda}^{0}$. To investigate this, we can incrementally modify $\bm{\lambda}^{0}$ by replacing its hyperparameter values with those from $\bm{\lambda}^{\ast}$, one at a time; a process known as ablation, widely used in empirical ML research (cohen-aimag1988; rendsburg-icml2020; herrmann-icml2024).
HPO-ablation studies were proposed by Fawcett-jh16a but limited to sequential single-hyperparameter ablation paths, ignoring interactions. Instead, we form an explanation game for ablation using all possible subsets, which allows us to capture interactions.
Definition 1 (Ablation Game).
The Ablation explanation game $\nu_{G_{A}}:2^{\mathcal{N}}\to\mathbb{R}$ is defined as a tuple $G_{A}:=(\bm{\lambda}^{0},\bm{\lambda}^{\ast},D,u),$ consisting of a baseline (default) configuration $\bm{\lambda}^{0}$, a target configuration $\bm{\lambda}^{\ast}$, a dataset $D$, and a measure $u$. Given a coalition $S\subseteq\mathcal{N}$, we construct an intermediate configuration with $\oplus_{S}:\bm{\Lambda}\times\bm{\Lambda}\to\bm{\Lambda}$ as
@@
and evaluate its value via $\nu_{G_{A}}(S):=\textsc{Val}_{u}(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0},D)\,\,.$
The Ablation game quantifies the worth of a coalition based on the comparison with a baseline configuration $\bm{\lambda}^{0}$. In XAI terminology, this approach is known as baseline imputation. Natural extensions of the Ablation game capture these ablations with respect to a distribution $\bm{\lambda}^{0}\sim p^{0}(\bm{\lambda}^{0})$ over configuration space $\bm{\Lambda}$ as $\mathbb{E}_{\bm{\lambda}^{0}\sim p^{0}(\bm{\lambda}^{0})}[\textsc{Val}_{u}(\bm{\lambda}^{\ast}\oplus_{S}\bm{\lambda}^{0},D)],$ which relates to the marginal performance (fANOVA). In XAI terminology, it is further distinguished between distributions $p(\bm{\lambda}^{0})$ that either depend (conditional) or do not depend (marginal) on the target configuration $\bm{\lambda}^{\ast}$. Baseline imputation is often chosen for efficiency and is also argued to have desirable properties (Sundararajan.2020b). Still, the choice of baseline strongly influences the explanation (sturmfels2020visualizing). We typically use a default configuration (AnastacioH20) here, though our methodology readily extends to probabilistic baselines.
5.2 Sensitivity and Tunability of Learners
Zooming out from a specific configuration, we can ask to what extent it is worthwhile to tune hyperparameters. In the literature, this question has been connected to the term of tunability (tunability). Tunability aims to quantify how much performance improvements can be obtained by tuning a learner, comparing against a baseline configuration, e.g., a configuration that is known to work well across various datasets (DBLP:conf/gecco/PushakH20a). In this context, we are interested in the importance of tuning specific hyperparameters. A classical tool to quantify variable importance is sensitivity analysis (Owen_2013), measuring the variance induced by the variables and decomposing their contributions into main and interaction effects.
Definition 2 (Sensitivity Game).
The Sensitivity game $\nu_{G_{V}}:2^{\mathcal{N}}\to\mathbb{R}$ is defined as a tuple $G_{V}:=(\bm{\lambda}^{0},\bm{\Lambda},p^{*},D,u),$ consisting of a baseline configuration $\bm{\lambda}^{0}$, a configuration space of interest $\bm{\Lambda}$ equipped with a probability distribution $p^{*}$, a dataset $D$, and a performance measure $u$. The value function is given by
@@ 5.2 Sensitivity and Tunability of Learners
A large value of a coalition $S\subseteq\mathcal{N}$ in the Sensitivity game indicates that these hyperparameters are important to be set to the right value. fANOVA implicitly rely on the Sensitivity game and compute the fANOVA decomposition, quantifying pure main and interaction effects. In game theory, this corresponds to the MIs of the Sensitivity game, which can be summarized using the SV and SIs (Fumagalli.2024a).
While sensitivity analysis is a suitable tool in XAI, it has some drawbacks for measuring tunability (tunability). First, as illustrated below, the total variance being decomposed $\nu_{G_{V}}(\mathcal{N})$ highly depends on the chosen probability distribution $p^{*}$ and the configuration space $\bm{\Lambda}$. Moreover, it does not reflect the performance increase expected when tuning all hyperparameters, but variations (in any direction). Second, for a coalition of hyperparameters $S\subseteq\mathcal{N}$, we expect that the coalition’s worth (performance) increases when tuning additional hyperparameters, i.e., $\nu(S)\leq\nu(T)$, if $S\subseteq T$. This property is known as monotonicity (Fujimoto.2006), but does not hold in general for the Sensitivity game $\nu_{G_{V}}$. For a simple example, we refer to the appendix. Based on tunability, we define an explanation game for tunability that exhibits monotonicity:
Definition 3 (Tunability Game).
The Tunability game is defined by a tuple $G_{T}=(\bm{\lambda}^{0},\bm{\Lambda},D,u),$ consisting of a baseline configuration $\bm{\lambda}^{0}\in\bm{\Lambda}$, a configuration space $\bm{\Lambda}$, a dataset $D$, and a measure $u$. The value function is given by
@@
The Tunability game directly measures the performance obtained from tuning the hyperparameters of a coalition $S$ while leaving the remaining hyperparameters at the default value $\bm{\lambda}^{0}$. The Tunability game is monotone, which yields the following lemma.
Proposition 1.
@@
While the main effects obtained from the Tunability game are non-negative, interactions can still be negative, indicating redundancies of the involved hyperparameters.
Comparing Tunability vs. Sensitivity.
We now showcase the different results of the Tunability game vs. the Sensitivity game using an educational example. We consider a two-dimensional configuration space $\bm{\Lambda}:=\Lambda_{1}\times\Lambda_{2}$ with discrete configurations $\Lambda_{1}:=\{0,1\}$ and $\Lambda_{2}:=\{0,\dots,m\}$ for $m>1$. The optimal configuration is defined as $\bm{\lambda}^{\ast}:=(1,m)$, and the performance is quantified by $\textsc{Val}_{u}(\bm{\lambda},D):=\mathbf{1}_{\lambda_{1}=\lambda_{1}^{\ast}}+\mathbf{1}_{\lambda_{2}=\lambda^{\ast}_{2}}$, where $\mathbf{1}$ is the indicator function. That is, we observe an increase of performance of $1$ for each of the hyperparameters set to the optimal configuration $\bm{\lambda}^{\ast}$. Lastly, we set the configuration baseline to $\bm{\lambda}^{0}:=(0,0)$ or $\bm{\lambda}^{0}:=\bm{\lambda}^{\ast}$. Intuitively, we expect that both hyperparameters obtain similar importance scores, since they both contribute equally to the optimal performance $\textsc{Val}_{u}(\bm{\lambda}^{\ast},D)=2$. Moreover, if the baseline is set to the optimal configuration $\bm{\lambda}^{\ast}$, we expect the score to reflect that there is no benefit of tuning. Since the hyperparameters affect the performance independently, we do not expect any interactions.
Table 1: Importance scores for a 2D HPO problem under the Sensitivity and Tunability games, with baseline set to $(0,0)$ and optimum $\bm{\lambda}^{\ast}$. Sensitivity assigns lower scores to hyperparameters with larger domains ($\lambda_{2})$. Setting $\bm{\lambda}^{0}=\bm{\lambda}^{*}$ reduces the Tunability scores to 0; Sensitivity is unaffected.
Game Sensitivity Tunability
@@
The HPI scores of the Sensitivity and Tunability game for the example are given by Table 1. Both approaches, Sensitivity and Tunability, correctly quantify the absence of interaction $\lambda_{1}\times\lambda_{2}$. As opposed to the Tunability game, the Sensitivity game assigns smaller scores to the hyperparameter $\lambda_{2}$ due to the larger domain $\Lambda_{2}$. In fact, the Sensitivity score of $\lambda_{2}$ roughly decreases with order $m^{-1}$. Moreover, the Tunability scores reflect the performance increase and, as expected, distribute the difference between the optimal and the baseline performance properly among the hyperparameters. In contrast, the Sensitivity scores decompose the overall variance, which depends on $\bm{\Lambda}$ and $p^{*}$. Lastly, setting the baseline configuration $\bm{\lambda}^{0}$ to $\bm{\lambda}^{\ast}$ decreases the Tunability scores to zero, whereas the Sensitivity scores remain unaffected. In summary, Sensitivity reflects the variability in performance when changing the hyperparameter values, whereas Tunability reflects the benefit of tuning over the baseline.
@@ 5.3 Optimizer Bias
The Tunability game aims to explain the importance of hyperparameters being tuned, which can also be used to gain insights into the capabilities of a hyperparameter optimizer. In particular, by comparing the optimal performance with the empirical performance of a single optimizer, we can uncover biases and pinpoint specific hyperparameters that the optimizer of interest fails to exploit. We define a hyperparameter optimizer as a function $\mathcal{O}:\mathbb{D}\times 2^{\bm{\Lambda}}\rightarrow\bm{\Lambda}$, mapping from the space of datasets and a configuration space to a specific configuration.
Definition 4 (Optimizer Bias Game).
The Optimizer Bias HPI game is defined as a tuple $G_{O}=(\bm{\Lambda},\bm{\lambda}^{0},\mathcal{O},D,u),$ consisting of a configuration space $\bm{\Lambda}$, a baseline $\bm{\lambda}^{0}$, the hyperparameter optimizer of interest $\mathcal{O}$, a dataset $D$ and a measure $u$. For $S\subseteq\mathcal{N}$, we define $\bm{\Lambda}^{S}:=\{\bm{\lambda}\oplus_{S}\bm{\lambda}^{0}:\bm{\lambda}\in\bm{\Lambda}\}$ and
@@
Intuitively, the value function captures how much performance is lost relative to the best known configuration. In other words, with the help of Definition 4, we can pinpoint where the hyperparameter optimizer $\mathcal{O}$ falls short, revealing, for example, whether it struggles to optimize certain hyperparameters or types thereof. The analysis can be conducted via inexpensive surrogate-based HPO benchmarks.
@@ 5.4 Practical Aspects of HyperSHAP
Naïvely, to evaluate a single coalition in Definition 3 of the Tunability game, we need to conduct one HPO run. While this can be costly, we argue that using surrogate models that are, e.g., obtained through Bayesian optimization, can be used to calculate the maximum efficiently. Surrogate models are commonly used in explainability methods for HPO, including fANOVA and related approaches (fANOVA; DBLP:conf/aaai/BiedenkappLEHFH17; DBLP:conf/nips/MoosbauerHCLB21; DBLP:conf/automl/SegelGTBL23). For HyperSHAP, we can bound the approximation error for the explanations as follows:
Theorem 1.
For a surrogate model with approximation error $\epsilon$, the approximation error of Shapley values and interactions in HyperSHAP is bounded by $2\epsilon$.
Optimizer Bias Analysis
To analyze Optimizer Bias, we propose to approximate $\nu_{G_{T}}$ using a diverse ensemble of optimizers $\mathbb{O}:=\{\mathcal{O}_{i}\}$, and choose the best result for $\bm{\Lambda}^{S}$ obtained through any optimizer from $\mathbb{O}$, forming a virtual optimizer, always returning the best-known value. This virtual best hyperparameter optimizer approximates
@@
Worst Case Analysis.
In order to identify hyperparameters that should not be mistuned, we can conduct a worst-case analysis with HyperSHAP by replacing the $\max$ by a $\min$ operator in Definition 3.
Game Extensions Across Multiple Datasets.
In a more general setting, we are interested in explanations across multiple datasets, for which we can extend the previous games naturally as follows:
Definition 5 (Multi-Dataset Games).
Given a collection of datasets $\mathcal{D}:=\{D_{1},\dots,D_{M}\}$, the corresponding games $\nu^{D_{i}}_{G}$ for $1\leq i\leq M$ with $G\in\{G_{A},G_{V},G_{T},G_{O}\}$, we define its multi-dataset variant with the value function $\nu^{\mathcal{D}}_{G}(S):=\bigoplus_{i=1}^{M}\nu^{D_{i}}_{G}(S)$, where $\bigoplus$ denotes an aggregation operator, e.g., the mean or a quantile of the game values obtained for the datasets $D_{i}$.
Considering explanations across datasets enables a broader view of the impact of how individual hyperparameters and their interactions affect generalization performance. Aggregating coalition values reveals which hyperparameters are generally worth tuning, rather than just data-specific importance, justifying tuning recommendations or uncovering systematic optimizer biases beyond data-specific effects.